# OmniDRCA: Parallel Speech-Text Foundation Model via Dual-Resolution Speech Representations and Contrastive Alignment

### v1

@@ 3.4 Multimodal Interaction Patterns
Practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Furthermore, Zhang et al. (2023a) found that the chain-of-modality (CoM) methodology can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in the Appendix. For training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
@@ 4 Experiments
Table 2: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks. Baseline results are from Yao et al. (2024) except LUCY (S2) and GLM-4-Voice-Small from Gao et al. (2025) and Zeng et al. (2025) respectively. Metrics are benchmark-specific (higher is better). The Avg. score averages four benchmarks (Gray: AlpacaEval excluded).
Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg.
GPT-4o-Realtime Closed-source 74 51.6 71.7 69.7 66.7
Step-Audio-Chat Text-Driven 130B 43 56.1 56.7 36.7 48.1
MiniCPM-o 2.6 Text-Driven 8B 51 40.0 61.0 40.2 48.1
Llama-Omni Text-Driven 8B 39 22.9 45.3 10.7 29.5
VITA-1.5 Text-Driven 7B 20 28.1 46.7 23.3 29.5
GLM-4-Voice Interleaved 9B 51 32.0 50.0 36.4 42.3
GLM-4-Voice-Small Interleaved 1.5B 5.4 18.3 4.6 9.4
LUCY (S2) Parallel 7B 25.6 54.0 22.8 34.1
Moshi Parallel 7B 24 23.8 43.7 16.7 27.1
Mini-Omni Parallel 0.5B 25 12.8 22.0 6.9 16.7
OmniDRCA Parallel 8B 45 30.0 65.0 32.9 43.2
OmniDRCA-Small Parallel 2B 38 22.1 55.3 17.9 33.3
Table 3: Speech-to-Text ($S\rightarrow T$) Generation Performance on OpenAudioBench using benchmark-specific metrics (higher is better). Baseline results are from KimiTeam et al. (2025), except LUCY (S2) and Moshi from Gao et al. (2025) and MiniCPM-o 2.6 from our experiments. Reasoning QA is removed since our model is only trained on English data. Avg. is averaged over the four benchmarks (Gray: AlpacaEval excluded).
Model Type Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
Qwen2-Audio-Chat Cascaded 7B 57.2 45.2 69.7 40.3 53.1
MiniCPM-o 2.6 Text-Driven 8B 64.1 69.2 78.0 63.0 68.5
Step-Audio-chat Text-Driven 130B 56.5 73.0 72.3 56.8 64.7
Baichuan-Omni-1.5 Interleaved 7B 59.6 58.7 74.3 55.4 62.0
GLM-4-Voice Interleaved 9B 57.9 55.4 76.0 51.8 60.3
Moshi Parallel 7B 26.6 62.3 22.8 37.2
LUCY (S2) Parallel 7B 29.3 57.3 23.3 36.6
OmniDRCA Parallel 8B 58.0 51.7 79.7 47.7 59.7
Table 4: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following (Zeng et al., 2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text. Baseline scores are copied from (Zeng et al., 2025), except MiniCPM-o 2.6 results are from our experiments.
@@
MiniCPM-o 2.6 Text-Driven 8B 4.18 13.31
Llama-Omni Text-Driven 8B 3.92 9.18
GLM-4-Voice Interleaved 9B 4.33 7.83
Moshi Parallel 7B 3.90 7.95
Mini-Omni Parallel 0.5B 3.17 25.28
OmniDRCA Parallel 8B 3.86 6.75
OmniDRCA-Small Parallel 2B 3.87 8.86
Table 5: Ablation study of Dual-Resolution Grouping Strategy Conducted on OmniDCRA-Small with Llama Questions Dataset.
Modal Group Size S2T T2M (T/S) T2T
S2T+T2T (Understanding) 1 55.33 61.33
5 62.33 69.67
T2M+T2T (Generation) 1 51.67 / 41.00 58.33
5 42.00 / 25.33 64.00
Table 6: Ablation study of Contrastive Cross-modal Alignment (Cont), SRM, and Multimodal Architecture Design (MD) on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
@@
OmniDRCA-Small 64.33 / 55.33 64.67 69.67 / 58.67 69.67 67.33 / 59.67 65.67 / 57.00 67.67 / 57.00
w/o Cont 61.67 / 53.00 62.33 70.00 / 60.00 74.00 69.33 / 61.00 63.00 / 55.00 66.33 / 58.67
   w/o SRM 21.67 / 15.33 56.00 45.22 / 35.00 73.00 64.33 / 50.67 55.67 / 42.33 40.33 / 27.67
   w/o MD 58.00 / 49.00 58.00 69.33 / 55.00 68.33
@@ 4.1 Experimental Setup
Datasets. Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for SRM pre-training. 17.8K hours of English speech conversations data synthesized from text are adopted for OmniDRCA training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^2] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b). Following prior works (Yao et al., 2024; Li et al., 2025), we evaluate the performance on the widely used Spoken Question Answering benchmarks, including AlpacaEval (Li et al., 2023), Web Questions (Berant et al., 2013), Llama Questions (Nachmani et al., 2024), and TriviaQA (Joshi et al., 2017). To ensure fair comparisons given that different systems employ distinct data subsets for evaluation, we adopt datasets from two established evaluation frameworks: UltraEval-Audio  [^3] https://github.com/OpenBMB/UltraEval-Audio for Speech-to-Speech ($S\rightarrow S$) evaluation and OpenAudioBench  [^4] https://huggingface.co/datasets/baichuan-inc/OpenAudioBench for Speech-To-Text ($S\rightarrow T$) evaluation. As discussed in Huang et al. (2025); Chen et al. (2025), the results on TriviaQA are for reference only and marked with “*” due to the absence of an official test set. Moreover, we detect performance anomalies of OmniDRCA on the UltraEval-Audio TriviaQA dataset; therefore, we replace UltraEval-Audio TriviaQA dataset with OpenAudioBench TriviaQA dataset for a more meaningful evaluation  [^5] For reference, this setup yields a score of 42.6 for MiniCPM-o 2.6 while its score on the original UltraEval-Audio TriviaQA is 40.2..
@@
Evaluation Metrics. Following the settings of UltraEval-Audio and OpenAudioBench, we utilize G-Eval (Liu et al., 2023) to evaluate model performance on AlpacaEval, and employ accuracy to assess model performance on other datasets. The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al. (2025).
@@
Baselines. Representive baselines are selected to cover diverse modeling paradigms: the closed-source model GPT-4o-Realtime (OpenAI, 2024b), cascaded system Qwen2-Audio-Chat (Chu et al., 2024), Text-Driven models including Step-Audio-Chat (Huang et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni (Fang et al., 2024), and VITA-1.5 (Fu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025) and GLM-4-Voice (Zeng et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Mini-Omni (Xie and Wu, 2024a). This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
@@
Implementation Details. SRM uses Qwen2.5-0.5B as its base LLM, trained following $T2M$ paradigm. OmniDRCA uses Qwen2.5-7B-Instruct as its base LLM, while OmniDRCA-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. The warmup rate is set to $1\%$ of the total training steps. For OmniDRCA, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for OmniDRCA-Small and SRM it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$. The AdamW method (Loshchilov and Hutter, 2019) is used for optimization. Experiments are run on 32x NVIDIA Tesla A100 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding. For OmniDRCA, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRM pretraining, and about $20$ hours on OmniDRCA post training.
@@ 4.2 Results and Discussions
Tables 2 and 3 compare OmniDRCA with the representative and competitive baselines across all major model categories, on speech-to-speech and speech-to-text generation performance. Table 2 shows that in prior works, best Text-Driven models substantially outperform top-tier Joint Speech-Text models, because text-driven systems do not allow speech tokens to interfere with the generation process, thereby minimally impacting the LLM’s generation paradigm. Moreover, prior interleaved models often outperform parallel models, though they exhibit higher latency (e.g., 13 text tokens for first speech in GLM-4-Voice) and lack temporal alignment, complicating multistream duplex modeling. As shown in Tables 2 and 3, OmniDRCA achieves SOTA performance under the parallel paradigm, achieving a substantial 24.9% relative gain over LUCY (S2) on the average score across three Speech-to-Speech QA benchmarks and 60.4% gain over Moshi across three Speech-to-Text QA benchmarks, excluding AlpacaEval in both cases. OmniDRCA also shows competitive performance against the interleaved GLM-4-Voice and Baichuan-Omni-1.5. However, a relatively large gap still remains when compared to text-driven systems. In future work, we plan to train the model on more extensive datasets to improve the performance.
The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (WER(%)), as shown in Table 4. Among models with comparable sizes, text-driven models tend to exhibit higher WER, while joint speech-text models achieve relatively lower WER. The model size also significantly impacts WER. OmniDRCA achieves the lowest WER, indicating strong alignment when generating both speech and text simultaneously. In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a). The gap in UTMOS between OmniDRCA and GLM-4-Voice suggests room for improvement in modeling prosody and preserving acoustic details.
@@
Dual-Resolution Grouping Strategy. Table 5 reveals the dual effects of our grouping strategy: while improving speech understanding (12.6% relative gain in S2T with group size 5) and text capabilities (13.6% relative gain in T2T), it degrades speech generation quality by 38.2% relative in T2M. On the other hand, as shown in Table 6, SRM demonstrates remarkable effectiveness in speech token generation, achieving 245.9% and 71.4% relative improvements in S2M and T2M tasks respectively. Notably, SRM also enhances speech comprehension metrics (S2T and STC-T) by 10-16% relative, suggesting that improved speech token reconstruction capability improves latent speech representations. The text-only performance (T2T) remains stable, confirming SRM’s modality-specific effectiveness without interfering with text processing pathways. Our dual-resolution architecture resolves this conflict by applying grouping only in the comprehension phase (5Hz) while using raw frame resolution (25Hz) with SRM for generation.
Contrastive Cross-modal Alignment. As shown in Table 6, the introduction of speech-text contrastive learning for sentence-level semantic alignment substantially enhances the model’s semantic understanding capabilities. While the final evaluation metric (S2S performance) shows improvement over the counterpart without contrastive loss, the enhancement exhibits modality-specific efficacy. For instance, S2T performance improves by 2.34% absolute (64.67 over 62.33), while T2T performance degrades by 4.33%, suggesting that while improving speech comprehension, the contrastive objective partially compromises the LLM’s native text understanding capacity. Our multimodal evaluation framework effectively reveals these nuanced impacts, guiding future work on mitigating text performance degradation through more balanced loss design.
@@
Multimodal Architecture Design (MD). Our multimodal design achieves two key goals: (1) context-aware modality selection and (2) implicit data augmentation. As shown in Table 6, on the speech generation task, STC/SAC/SUC demonstrate substantial improvements over S2M. STC achieves 61.00, substantially surpassing the direct generation baseline of 55.33. This reveals the critical role of CoM while also indicating potential optimization opportunities for our S2M model. Without incorporating the three CoM variants (STC/SAC/SUC), the model fails to adopt this generation paradigm guided by system prompts. Final S2M performance shows that training without MD results in a 7.5% performance degradation, confirming the data augmentation effect of our approach.
Inspired by Wang et al. (2025), to enable full-duplex interaction, we propose a time-division multiplexing (TDM) input stream that allows OmniDRCA to accept user speech inputs during the assistant’s speech generation phase, thereby efficiently utilizing otherwise idle time slots. The resulting OmniDRCA-Duplex achieves full-duplex dialogue capabilities. The methodology and empirical validations of OmniDRCA-Duplex are detailed in Appendix.
5 Conclusions, Limitations, and Future Work
We introduce OmniDRCA, a novel parallel speech-text foundation model that integrates joint autoregressive modeling with dual-resolution speech representations and contrastive cross-modal alignment. Experiments show that OmniDRCA sets new SOTA speech-to-speech/text generation performance among parallel models and reaches competitive performance to interleaved models.
Limitations: (1) OmniDRCA’s speech generation quality remains constrained by suboptimal speech tokenization. We will explore other tokenization methods, such as CosyVoice’s speech tokenizer, to address this limitation. (2) We will extend exploration from current speech understanding and generation tasks to broader audio comprehension/synthesis tasks, and integrate visual modality.
Appendix A OmniDRCA-Duplex
As mentioned in Introduction, we explore the potential of OmniDRCA for full-duplex voice conversation. By leveraging time-division multiplexing, we introduce a novel spoken dialogue model, OmniDRCA-Duplex, which supports real-time responses to user interruptions with minimal performance degradation. In this section, we first introduce OmniDRCA-Duplex in Appendix A.1, then present the experimental results and discussions of its full-duplex conversation capabilities in Appendix A.2.
A.1 Methodology
As shown in Figure 2, compared with OmniDRCA, the overall architecture of OmniDRCA-Duplex remains largely unchanged, which is also composed of the multimodal LLM (MLLM) and the speech refinement model (SRM). The primary differences between OmniDRCA-Duplex and the original OmniDRCA model lie in two aspects. The first aspect is the organization of the input structure. OmniDRCA-Duplex organizes user speech and assistant speech alternately along the temporal axis as input, referred to as the Time-Division Multiplexing Input Stream, while OmniDRCA adopts turn-based speech as the same as text-based LLM. For the second aspect, to enhance the model’s sensitivity to user interruptions, an additional assistant state prediction head is introduced in OmniDRCA-Duplex to determine whether muting should be activated. The following paragraphs provide detailed descriptions of these two designs.
Figure 2: The architecture of OmniDRCA-Duplex.
Time-Division Multiplexing Input Stream.
Full-duplex spoken dialogue models allow simultaneous receptions of inputs from both user and assistant. Inspired by Wang et al. (2025), OmniDRCA-Duplex organizes inputs into a double-channel format, specifically, an alternating sequence of user speech and assistant speech along the temporal axis. Besides, same position id is used in one time slot, which paired user and assistant speech representations together. The model enforces attention masking between concurrent speech representations in individual time slots since they appear simultaneously. This structured arrangement, termed time-division multiplexing (TDM) input stream, is visually illustrated in Figure 2 TDM-Duplex. This approach allows the model to accept user speech when the assistant is generating speech, effectively utilizing the time slots that would otherwise be idle. The TDM input stream is designed to handle both user and assistant speech inputs, ensuring a seamless interaction experience.
Awareness of Interruptions.
As our early experiments show, the model may not be aware of user interruptions during the generation of assistant speech, since the model prefers to generate speech based on the content in the time slots from its own role (that is, Assistant Speech Token Stream). This can lead to a lack of responsiveness to user interruptions and a less engaging interaction experience. To address this issue, we implement a mechanism that enhances the model’s awareness to user interruptions during assistant speech generation. We achieve this goal by adding an additional head to predict the assistant’s speech generation state. The model is trained to predict whether the assistant is currently generating speech or not, based on the input speech tokens. This additional head provides the model with contextual information about the assistant’s speech generation state, enabling it to respond more effectively to user interruptions.
A.2 Experiments
Experimental Setup.
The original speech datasets using for OmniDRCA post-training are edited to simulate full-duplex conversation scenarios. Two types of datasets are constructed: 1) Non-interruption Dataset: Dialogue turns are structured into dual-channel audio. To better simulate human interactions, a response time of $0\sim 400$ ms is randomly inserted between dialogue turns. This dataset enables the model to learn when to initiate and terminate speech. 2) Interruption Multi-Turn Dialogue Dataset: The assistant’s responses in dialogue turns are randomly truncated, followed by inserting a new user input , before constructing dual-channel datasets. Through this data simulation approach, the model learns to handle incomplete interrupted semantics. In total, about 4K hours non-interruption dataset and 4K hours interruption multi-turn dialogue dataset are adopted for OmniDRCA-Duplex training. UltraEval-Audio Benchmark is used for evaluating the semantic ability, while multi-turn Llama Questions randomly pairing two of questions together is used for evaluating the interruption ablitily. Specific, providing these two rounds Llama Questions examples, if the model can mute itself while the user is speaking and respond after the user finishes, it is considered a successful Turn-taking. The testing evaluates the success rate of Turn-taking. 666 Here, Turn-taking@k Metrics is not conducted unlike other duplex speech models (Zhang et al., 2024a). This approach stems from intentional dataset construction aimed at endowing the model with human-like response timing and enabling scenarios where semantic information coexists simultaneously in both channels. The design specifically emphasizes making the model’s turn-taking response time more flexible rather than prioritizing the speed of turn-taking as the objective for the duplex model.
Table 7: Speech-to-Speech ($S\rightarrow S$) Generation Performance on UltraEval-Audio Spoken Question Answering Benchmarks Comparing the original OmniDRCA model and its Duplex Model.
Model Size AlpacaEval Web Q. Llama Q. TriviaQA Avg
Moshi 7B 24 23.8 43.7 16.7 27.1
OmniDRCA-Duplex-Small 2B 33 16.1 46.6 13.6 27.3
OmniDRCA-Small 2B 38 22.1 55.3 17.9 33.3
Results.
As demonstrated in Table 7, evaluation results on the UltraEval-Audio Spoken Question Answering Benchmark indicate that OmniDRCA-Duplex-Small attains an average score of 27.3. This represents a 18.2% relative performance degradation compared to the baseline OmniDRCA-Small model’s average score of 33.3. Notably, the experimental results show that OmniDRCA-Small-Duplex (2B model size) has a slight performance advantage over the larger 7B full-duplex model Moshi (27.3 versus 27.1). For the model’s interruption performance, by setting the multi-turn Llama Questions with intervals of $t=3s$ between turns, OmniDRCA-Small-Duplex achieves 98.64% success rate of Turn-taking, meaning the model can keep silent when user is talking and consistently respond to user-interrupted utterances in the second round.
Table 8: The Impact of SRM Pretraining Data on Llama Questions Performance.
Model S2M (T/S) S2T T2M (T/S) T2T
OmniDRCA-Small (Less) 52.67 / 44.33 60.33 67.67 / 53.00 75.00
   w/o. SRM-PT 38.33 / 30.33 56.00 59.33 / 46.33 73.33
Figure 3: Performance Scaling of OmniDRCA-Small (w/o. Cont) on LLaMA Question Benchmark.
Appendix B Dataset Scaling
Pre-Training
To examine the impact of pretraining data on model performance, the SRM pretraining step is eliminated, and the OmniDRCA-Small model is retrained. For computational efficiency, fewer post-training steps are adopted, called OmniDRCA-Small (Less), leading to the inferior performance compared to the results of OmniDRCA-Small reported in the main text. As shown in Table 8, SRM pretraining has the most significant impact on the performance of S2M and T2M, while its effect on S2T and T2T is relatively minor. This is reasonable because SRM is primarily responsible for refining speech token generation tasks, thus exerting a smaller influence on purely textual generation. This underscores the importance of SRM pretraining.
Post-Training

### v2

@@ 3.3 CoM-Mixing Training strategy
The chain-of-modality (CoM) methodology (Zhang et al., 2023a) can enhance performance by leveraging intermediate textual transcriptions, which is particularly suitable for scenarios where real-time processing is not critical. Furthermore, practical applications involve scenarios where both speech and text output or text-only output are required, necessitating the system to dynamically generate different modalities based on specific needs. Seven interaction patterns for diverse input-output requirements are summarized in Table 1. System prompts are employed to regulate the model’s output behavior. For example, prompts such as ‘‘[System] You are a helpful assistant and asked to generate both text and speech tokens at the same time.’’ explicitly guide the model to produce multimodal outputs. Detailed prompts are shown in the Appendix A. During inference, specifying these system prompts enables the generation of desired output results. For CoM-Mixing training, we construct data variants following the seven interaction patterns and obtain a mixture of data for training the model. During inference, user can manually prepend the appropriate system prompt to the input sequence to meet the output requirement. This flexibility ensures adaptability to varying modality generation demands.
@@ 4 Experiments
Table 2: Speech-to-Text (S $\rightarrow$ T) and Speech-to-Speech (S $\rightarrow$ S) generation performance on spokenQA benchmarks. The results for the S2T baselines are taken from the previous SOTA model Kimi-Audio’s paper (Xu et al., 2025), and the S2S results are taken from the previous SOTA model VITA-Audio’s paper (Long et al., 2025). Otherwise, results are taken from their original papers. The model size refers only to the parameter count of the LLM backbone, excluding the audio encoder, audio tokenzier, and audio detokenzier module. Frame Rate (In/Out) denotes the input speech frame rate and the output speech plus text frame rate for the LLM backbone. The average number of text tokens corresponding to one second of speech is denoted by $\tau$. Metrics are benchmark-specific (higher is better). The Avg. score averages two benchmarks. The best results in each group are highlighted in bold, while the second-best results are underlined.
Model Size Frame Rate (In/Out) Llama Q. Web Q. Avg.
$\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$ $\rightarrowT S\rightarrowS$
@@
DrVoice-Small (Our) 2B 5/5 73.0 56.0 42.1 31.1 57.5 43.5
Moshi (2024) 7B 12.5/12.5 62.3 21.0 26.6 9.2 44.5 15.1
[…]
LUCY (S2) (2025) 7B 12.5/48 57.3 51.0 29.3 18.2 43.3 34.6
[…]
Kimi-Audio (2025) 8B 12.5/12.5 79.3 - 70.0 - 74.7 -
[…]
Stream-Omni (2025) 7B 25/25 76.3 65.0 44.2 27.5 60.3 46.3
[…]
DrVoice (Our) 8B 5/5 85.0 71.0 65.2 40.1 75.1 55.5
Table 3: Speech quality performance on the collections of Spoken Quesion Answering Benchmarks following Zeng et al. (2025), in terms of UTMOS for assessing overall speech quality and ASR WER for assessing alignment between generated speech and text. Baseline scores are copied from Zeng et al. (2025). FR(In/Out) indicates the input speech frame rate and the output (speech + text) frame rate for the LLM backbone, while $\tau$ represents the average text tokens per second of speech. The best results in each group are bolded.
[…]
Mini-Omni (2024a) 0.5B 50/48 3.17 25.28
DrVoice-Small 2B 5/5 3.87 8.86
[…]
Moshi (2025) 7B 12.5/12.5 3.90 7.95
[…]
DrVoice 8B 5/5 3.86 6.75
Table 4: Impact of the Grouping Factor on Llama Q. S2T and S2M performance. Experiments were conducted on DrVoice-Small w/o. CSE, using data subsets for faster exploration. The S2T model was trained exclusively on mixed S2T and T2T data, while the S2M model was trained on mixed S2M and T2T data.
Grouping Factor S2T S2M (T/S)
1 55.67 4.00 / 2.67
3 64.67 15.67 / 5.00
5 63.33 37.67 / 28.00
7 62.67 36.00 / 16.67
Table 5: Ablation study of Continuous Speech Encoder (CSE), Speech Refined Head (SRH), SRH-Pretraining and CoM-Mixing Training on the Llama Questions dataset. S2M, S2T, etc are defined in Table 1.
@@
DrVoice-Small 68.67 / 56.00 72.33 72.33 / 56.00 75.33 75.67 / 68.33 71.67 / 62.67 73.33 / 62.00
w/o. CSE 61.67 / 53.00 62.33 70.00 / 60.00 74.00 69.33 / 61.00 63.00 / 55.00 66.33 / 58.67
   w/o. SRH-Pretraining 38.33 / 30.33 56.00 59.33 / 46.33 73.33 67.33 / 57.67 54.00 / 42.33 54.33 / 42.67
    w/o. SRH 21.67 / 15.33 56.00 45.22 / 35.00 73.00 64.33 / 50.67 55.67 / 42.33 40.33 / 27.67
   w/o. CoM-Mixing 58.00 / 49.00 58.00 69.33 / 55.00 68.33 –/– –/– –/–
Figure 2: Computational Resources across different Grouping Factor.
Figure 3: Performance Scaling of DrVoice-Small (w/o. Continuous Speech Encoder) on LLaMA Question Benchmark.
@@ 4.1 Experimental Setup
Approximately 210K hours of audio-text paired data for speech-text modality synchronization is adopted for pre-training speech refined head. 17.8K hours of English speech conversations data synthesized from text are adopted for DrVoice training: (1) VoiceAssistant (Xie and Wu, 2024a, b), which contains human-agent interactions; (2) Spoken Alpaca  [^1] https://huggingface.co/datasets/GSQA/spoken-alpaca-gpt4, which provides instructed dialogues; and (3) Synthesized Datasets, which are synthesized from text datasets such as Infinity-Instruct (BAAI, 2024), using the CosyVoice TTS system (Du et al., 2024a, b). Following the common practice in prior works (Long et al., 2025; KimiTeam et al., 2025; Yao et al., 2024; Zhang et al., 2025), the Speech-to-Speech ($S\rightarrow S$) and Speech-To-Text ($S\rightarrow T$) performance is evaluated on widely used Spoken Question Answering benchmarks, including Web Questions (Berant et al., 2013) and Llama Questions (Nachmani et al., 2024).
Evaluation Metrics.
@@
Accuracy is emplyed to assess model performance on datasets. The generated speech is transcribed to text using whisper-v3-large model (Radford et al., 2022), then word error rate (WER) of transcripts is computed against the generated text to assess the alignment between generated speech and text, as well as using UTMOS (Saeki et al., 2022) to evaluate the overall speech quality, following Zeng et al. (2025).
Baselines.
@@
Representive baselines are selected to cover diverse modeling paradigms: the cascaded system Audio Flamingo 3 (Zhang et al., 2025), Text-Driven models including Freeze-Omni (Défossez et al., 2024), MinMo (Chen et al., 2025), MiniCPM-o 2.6 (Yao et al., 2024), Llama-Omni2 (Fang et al., 2025), and Qwen2.5-Omni (Xu et al., 2025), and among Joint Speech-Text models, interleaved models including Baichuan-Omni-1.5 (Li et al., 2025), GLM-4-Voice (Zeng et al., 2025), Stream-Omni (Zhang et al., 2025), and VITA-Audio (Long et al., 2025), alongside parallel models including LUCY (Gao et al., 2025), Moshi (Défossez et al., 2024), and Kimi-Audio (Xu et al., 2025). Noting that Kimi-Audio achieves state-of-the-art (SOTA) performance on Speech-to-Text Question Answering, while VITA-Audio sets the SOTA on Speech-to-Speech Question Answering. This configuration enables systematic comparisons across mainstream speech-text modeling strategies.
Implementation Details.
@@
Qwen2.5-0.5B is trained following $T2M$ paradigm with Speech-text Alignment data to initialize SRH (SRH-PT). DrVoice uses Qwen2.5-7B-Instruct as its base LLM, while DrVoice-Small utilizes Qwen2.5-1.5B-Instruct as the base LLM. Grouping factor is set to $k=5$. The warmup rate is set to $1\%$ of the total training steps. For DrVoice, the max learning rate is set to $1\times 10^{-5}$ and then decayed via cosine annealing to $1\times 10^{-6}$, while for DrVoice-Small and SRH-PT it is set to $1\times 10^{-4}$ and then decayed to $1\times 10^{-5}$. The AdamW method (Loshchilov and Hutter, 2019) is used for optimization. Experiments are run on 32x NVIDIA Tesla A800 80G GPUs with Brain floating-point format (BF16(Kalamkar et al., 2019) to accelerate training and decoding. For DrVoice, DeepSpeed ZeRO-2 (Rajbhandari et al., 2020) is implemented to prevent GPU out-of-memory. It takes about $20$ hours on SRH-PT, and about $45$ hours on DrVoice post training.
@@ 4.2 Results and Discussions
[…]
The quality of generated speech is evaluated in terms of overall speech quality (UTMOS) and alignment between generated speech and text (ASR-WER(%)), as shown in Table 3. Among models with comparable sizes, DrVoice achieves low ASR-WER, indicating strong alignment when generating both speech and text simultaneously. In contrast, UTMOS scores are more constrained by the speech tokenizer and detokenizer (Hassid et al., 2023; Du et al., 2024a). The gap in UTMOS between DrVoice and GLM-4-Voice suggests room for improvement in speech naturalness and intelligibility.
@@
Continuous Speech Encoder.
As shown in Table 5, the Continuous Speech Encoder (CSE) is vital for tasks involving speech inputs. Removing it from DrVoice-Small leads to a significant performance drop in speech understanding and speech generation. Specifically, the S2T score decreases from 72.33 to 62.33 (13.8% relatively), and the S2M (T) score falls from 68.67 to 61.67 (10.2% relatively). In contrast, the impact on text-only tasks is minimal, with the T2T score only slightly decreasing from 75.33 to 74.00. This confirms the CSE’s effectiveness in processing and representing speech information.
Dual-Resolution.
Our dual-resolution approach combines input grouping for efficiency and comprehension with a refinement head for high-quality generation.
Grouping Factor. Table 4 and Figure 2 demonstrate that our grouping strategy significantly enhances both performance and computational efficiency. Contrary to degrading generation, grouping substantially improves both speech understanding (S2T) and speech-to-speech generation (S2M). For instance, increasing the grouping factor from 1 to 5 raises the S2T score from 55.67 to 63.33 (a 13.7% relative gain). The S2M (T/S) score sees a significant improvement, peaking at 37.67 / 28.00 with a grouping factor of 5. Furthermore, as shown in Figure 2, using a grouping factor of 5 instead of 1 reduces nearly 50% GPU hours in each setting, suggesting the efficiency of grouping machinism.
Speech Refinement. As shown in Table 5, the Speech Refinement Head (SRH) is remarkably effective for speech token generation tasks. By comparing the model with SRH (w/o. SRH-Pretraining) to one without (w/o. SRH), we observe that adding SRH achieves a 76.9% relative improvement in S2M (T) (from 21.67 to 38.33) and a 31.2% relative improvement in T2M (T) (from 45.22 to 59.33). The text-only performance (T2T) remains stable, confirming that SRH enhances speech generation without interfering with text processing pathways. Our dual-resolution architecture effectively combines these components, using grouping for efficient comprehension and SRH at the raw frame resolution for speech generation.
To examine the impact of pretraining, we remove the SRH pretraining step and retrain the model. As shown in Table 5, removing pretraining (comparing w/o. CSE with w/o. SRH-Pretraining) has the most significant impact on speech generation, causing S2M (T) performance to drop by 37.8% and T2M (T) by 15.2%. The effect on S2T is smaller (10.2% drop), and it is negligible for T2T. This underscores the critical importance of SRH pretraining for refining speech token generation.
CoM-Mixing Training Strategy.
@@
As shown in Table 5, where tasks guided by contextual system prompts (STC/SAC/SUC) demonstrate substantial improvements over direct S2M generation. For example, STC (T) achieves a score of 75.67, significantly surpassing the S2M (T) baseline of 68.67. This shows the model successfully learns to adopt the generation paradigm guided by system prompts. The data augmentation effect is confirmed by ablating the CoM-Mixing strategy entirely. Training without it (w/o. CoM-Mixing Training) results in a 15.5% relative performance degradation in S2M (T) (from 68.67 to 58.00), confirming the value of our mixed-modality training approach.
Data Scaling